How much heat, in joules, is absorbed by 500.0 g of water as its temperature is increased from 10.0oC to 85.0oC?  The specific heat of water is 4.184 J/g•oC 



1
Expert's answer
2021-05-10T01:50:16-0400

We know the formula,


"Q= nC_v \Delta T"


"Q = \dfrac{500}{18} \times 4.184 \times (85-10)"


"Q = 8716.66J"